OpenStudy (anonymous):

Last year, Noah purchased 64 shares of Stock A at $51 per share, 84 shares of Stock B at $85 per share, and a four-year $3500 bond with an 11.83% coupon for $3550. Noah sold both stocks today. Stock A is worth $58 per share and Stock B has a value of $90 per share. Assuming neither stock paid a dividend, which investment has the highest rate of return? Stock A Stock B Bond Stock A and Stock B

OpenStudy (anonymous):

1ddddd !

OpenStudy (anonymous):

haha :P

jimthompson5910 (jim_thompson5910):

According to this page http://www.investopedia.com/university/bonds/bonds3.asp it says that bond yield is calculated by the formula Yield = coupon/price In this case, the bond has a coupon of 0.1183*3500 = 414.05 So the yield for the bond is Yield = coupon/price = 414.05/3550 = 0.1166338028169 which is roughly 11.66%

jimthompson5910 (jim_thompson5910):

what rates of returns did you get for stock A and stock B?

OpenStudy (anonymous):

i can't remember the formula D:

jimthompson5910 (jim_thompson5910):

for stock A, what is the initial price and final price?

OpenStudy (anonymous):

The rate of return for Stock A: 13.72% for Stock b: 5.88% i think

OpenStudy (anonymous):

is that right?

jimthompson5910 (jim_thompson5910):

very good

jimthompson5910 (jim_thompson5910):

so stock A is the winner

OpenStudy (anonymous):

so a?

OpenStudy (anonymous):

I have one more question and i promise i won't bother you again tonight!